OpenStudy (anonymous):

find the volume under the graph of the function f(x,y)=6x^2y over the region of the triangle with verticies of (1,3)(3,9)(3,3)

OpenStudy (anonymous):

fun double integral problem

OpenStudy (anonymous):

(1,3)(3,9)(3,3) derive function

OpenStudy (anonymous):

right..i'm there

OpenStudy (anonymous):

i get y = 3x and y = 3....and the third one is undefined

OpenStudy (anonymous):

but this is where i'm lost in setting up the integral i guess

OpenStudy (anonymous):

should be the integral of 1 to 9 and the other integral for terms of x is screwing with me

OpenStudy (anonymous):

sorry 3 to 9

OpenStudy (anonymous):

choose dx dy x should go from x=1/3y to x=3 y should go from 3 to 9

OpenStudy (anonymous):

can u explain how u got the x values?

OpenStudy (anonymous):

i'm confused

OpenStudy (anonymous):

right...the x = 1/3y...how so?

OpenStudy (anonymous):

since the slope is 0....y = 3

OpenStudy (anonymous):

so is it just x = 1/3y?

OpenStudy (anonymous):

and an assumed x value?

OpenStudy (anonymous):

y=3x (as you found out) solve for x y/3= x

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

ur using the same one okay

OpenStudy (anonymous):

ty sooo much

OpenStudy (anonymous):

\[\int _3^9\int _{\frac{y}{3}}^31dxdy=\int _1^3\int _3^{3x}1dydx\]

OpenStudy (zarkon):

what's more important is that \[\int _3^9\int _{\frac{y}{3}}^36x^2y\,dxdy=\int _1^3\int _3^{3x}6x^2y\,dydx\]